Independent random samples of professional football and basketball players gave the following information. Assume that the weight distributions are mound-shaped and symmetric.

Weights (in lb) of pro football players: x1; n1 = 21

244 262 255 251 244 276 240 265 257 252 282
256 250 264 270 275 245 275 253 265 271

Weights (in lb) of pro basketball players: x2; n2 = 19

205 200 220 210 192 215 222 216 228 207
225 208 195 191 207 196 182 193 201

(a) Use a calculator with mean and standard deviation keys to calculate x1, s1, x2, and s2. (Round your answers to one decimal place.)

x1 =
s1 =
x2 =
s2 =


(b) Let μ1 be the population mean for x1 and let μ2 be the population mean for x2. Find a 99% confidence interval for μ1μ2. (Round your answers to one decimal place.)

lower limit    
upper limit    
Answer #1

Cuiven theet 259 62 Sj2冫147.95 2 s2-16383 99 cant idence intes va tos population mear 14 一20% 147 95+18 χ 163.83 214 ig- 2 2loues limit 43.3
Answer #2
football ( X ) Σ ( Xi- X̅ )2 basketball ( Y ) Σ ( Yi- Y̅ )2
246 190.44 202 13.69
263 10.24 200 32.49
254 33.64 220 204.49
251 77.44 210 18.49
244 249.64 192 187.69
276 262.44 215 86.49
240 392.04 221 234.09
265 27.04 216 106.09
257 7.84 228 497.29
252 60.84 207 1.69
282 492.84 225 372.49
256 14.44 208 5.29
250 96.04 195 114.49
264 17.64 191 216.09
270 104.04 207 1.69
275 231.04 196 94.09
245 219.04 182 561.69
275 231.04 193 161.29
253 46.24 201 22.09
265 27.04
272 148.84
Total 5455 2939.84 3909 2931.71

Mean X̅ = Σ Xi / n
X̅ = 5455 / 21 = 259.8
Sample Standard deviation SX = √ ( (Xi - X̅ )2 / n - 1 )
SX = √ ( 2939.84 / 21 -1 ) = 12.1

Mean Y̅ = ΣYi / n
Y̅ = 3909 / 19 = 205.7
Sample Standard deviation SY = √ ( (Yi - Y̅ )2 / n - 1 )
SY = √ ( 2931.71 / 19 -1) = 12.8

part a)

X1 = 259.8

S1 = 12.1

X2 = 205.7

S2 = 12.8

Part b)

Confidence interval :-
(X1 – X2) + ta/2,DF) (S}/n1) + (S}/n2)
t(α/2, DF) = t(0.01 /2, 37 ) = 2.715

DF = ((S/n1+ s/n2))/((S/nl)/n1 - 1) + (S/n2)/n2 - 1))
DF = ((12.124/21+12.7621-/19))/((12.1242/21-/21-1)+(12.7621-/19)/19-
DF = 37

(259.7619 - 205.7368) + 10.01/2.37 (12.1242/21) + (12.76212/19)
Lower Limit = (259.7619 - 205.7368) - €0.01/2.37 (12.1242/21) + (12.76212/19)
Lower Limit = 43.3
Upper Limit = (259.7619 - 205.7368) + +0.01/2.37 (12.1242/21) + (12.76212/19)
Upper Limit = 64.7
99% Confidence interval is ( 43.3 , 64.7 )